OpenStudy (anonymous):

Hi, can you explain why x^-1/2 = 1/2√x? Thanks.

OpenStudy (anonymous):

\[x^{\frac{1}{2}}=\sqrt{x}\] it is just exponential notation

OpenStudy (anonymous):

the minus sign means take the reciprocal, so \[x^{-\frac{1}{2}}=\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

there is no two in it though, so maybe the question was about something else

OpenStudy (anonymous):

Oops I meant why \[\frac{ 1 }{ 2 }x ^{-\frac{ 1 }{ 2 }} = \frac{ 1 }{ 2\sqrt{x}}\]

OpenStudy (anonymous):

same reason

OpenStudy (anonymous):

Thank you :))

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

let me guess, derivative?

OpenStudy (anonymous):

Yes!